OIL COMPANY is planning to issue new shares in the stock market and the management contacted with various investment banks. In this regard, NBK Investment bank offer KUWAIT OIL COMPANY following options for its IPO sales of equity;


a) A best effort arrangement where NBK will keep $2 per share for each share sold
b) A firm commitment arrangement of $20,000,000
KUWAIT OIL COMPANY plans to issue 2000,000 shares at $11 per share to public, please analyze;
i. In case only 100% shares are sold which, is better choice for KUWAIT OIL COMPANY; which is better choice for NBK Investment Bank?
ii. In case only 90% shares are sold which, is better choice for KUWAIT OIL COMPANY which is better choice for NBK Investment Bank?
iii. In case only 75% shares are sold which, is better choice for KUWAIT OIL COMPANY; which is better choice for NBK Investment Bank?
Answer #1

Answer i.

Analysing better choice for Kuwait Oil Company

Cash Inflows in option a (2000,000*11)-(2,000,000*2) =18,000,000

Cash Inflows in option b = 20,000,000

Hence firm commitment arrangement is better choice for Kuwait Oil company

Analysing better choice for NBK

Cash Inflows in option a (2,000,000*2) =4,000,000

Cash Inflows in option b = (22,000,000-20,000,000)= 2,000,000

Hence, best effort arrangement is better choice for NBK

Answer ii

Analysing better choice for Kuwait Oil Company

Cash Inflows in option a (1,800,000*11)-(1,800,000*2) =16,200,000

Cash Inflows in option b = 20,000,000

Hence firm commitment arrangement is better choice for Kuwait Oil company

Analysing better choice for NBK

Cash Inflows in option a (1,800,000*2) =3,600,000

Cash Inflows in option b = (19,800,000-20,000,000)= (200,000)

Hence, best effort arrangement is better choice for NBK

Answer iii

Analysing better choice for Kuwait Oil Company

Cash Inflows in option a (1,500,000*11)-(1,500,000*2) =13,500,000

Cash Inflows in option b = 20,000,000

Hence firm commitment arrangement is better choice for Kuwait Oil company

Analysing better choice for NBK

Cash Inflows in option a (1,500,000*2) =3,000,000

Cash Inflows in option b = (16,500,000-20,000,000)= (3,500,000)

Hence, best effort arrangement is better choice for NBK